OpenStudy (anonymous):

Find sin2x and cos2x when tanx= sqrt 2. PLEASE HELP SOLVE

OpenStudy (anonymous):

You mean sin^2(x) or sin(2x)?

OpenStudy (anonymous):

sin2x so double angle formula.. not sure where to begin with tangent

OpenStudy (anonymous):

Use pythagorus to find ? and go from there

OpenStudy (anonymous):

ok that was wrong!

OpenStudy (anonymous):



OpenStudy (anonymous):

so you have \[\sin(x)=\frac{\sqrt{2}}{\sqrt{3}}\] and \[\cos(x)=\frac{1}{\sqrt{3}}\] use those to find \(\sin(2x)\) and \(\cos(2x)\)

OpenStudy (anonymous):

how did you find rad 3 as the hypotenuse?

OpenStudy (anonymous):

they hypotenuse is \[\sqrt{1^2+\sqrt{2}^2}=\sqrt{1+2}=\sqrt{3}\] by pythagoras

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

okay with that?

OpenStudy (anonymous):

yeah that clicks to me now! thank you so much

OpenStudy (anonymous):

yw it is good from there, right? \[\sin(2x)=2\sin(x)\cos(x)=2\times \frac{1}{\sqrt3}\times \frac{\sqrt2}{\sqrt3}=\frac{2\sqrt2}{3}\] etc

OpenStudy (anonymous):

got it from there!! thanks again